# Multi-intention Inverse Q-learning for Interpretable Behavior Representation

### v3

@@ 1 Introduction
Characterizing decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience [Niv09, WC19]. Prior research has formulated a variety of mathematical behavioral models across diverse tasks [ARS+22, BNLS22], including generalized linear models and models based on reinforcement learning. These forward models facilitate the understanding and comparison of decision-making strategies employed by both human and animal subjects. Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities [HDB+19, HTAW22]. Forward models require an empirically defined reward function that guides subjects optimizing their behavior during decision-making. However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks [AKK+23, RZPM21]. Inverse reinforcement learning (IRL) [NR+00, AD21] is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was maximizing its long-term return. Along with the significant successes of IRL in autonomous driving [KHWB20, NSA23], robotics [KZH+23, CLA23], and healthcare domains [CNDPP20, CvdS21], it appears to be emerging as a valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by [YNI+18, KDSP20, AKK+23].
@@
Classic IRL methods seek to identify a single, fixed reward function for a specific scenario. In contrast, Ashwood et al. [AJP22] suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity. Under this assumption, they proposed the dynamic inverse reinforcement learning (DIRL) framework, which parametrizes the animal’s reward function as a smoothly time-varying linear combination of a small number of spatial reward maps, which are referred to as ‘goal maps’. By assuming the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time. This innovative framework achieved state-of-the-art performance in animal behavior prediction. Nevertheless, demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by [ARS+22] that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
@@
To address this requirement, we propose the novel class of hierarchical inverse Q-learning (HIQL) algorithms, which extend the fixed-reward inverse Q-learning (IQL) framework from [KHWB20] to solve multi-intention IRL problems. Based on the assumption that the intention transition dynamics follows a Markov process, HIQL integrates an expectation-maximization (EM) approach to first divide expert trajectories into multiple intention segments in an unsupervised manner, and then solve the IRL problem independently for each segment. We then compare the performance of HIQL and the state-of-the-art model DIRL [AJP22] in a simulated gridworld environment and a dataset of trajectories from real mice navigating a $127$-node-labyrinth [RZPM21]. The results show that our HIQL algorithm outperforms DIRL in behavior prediction in both benchmarks, and produces interpretable reward functions. Finally, we applied HIQL in a real mice decision-making dataset from a dynamic two-armed bandit task [DLCSS+23], and mathematically characterized exploitation and exploration behavior of animals during value-based decision-making.
@@ 2 Related work
Various approaches have been introduced to address multi-intention IRL problems. Notably, several frameworks based on parametric [BMSL11, LMR+21], or non-parametric [CK12, BMJD21] approaches allow for learning from multiple agents with distinct reward functions. Based on the assumption that the expert sticks to the same intention through one episode, these methods iterate between clustering expert demonstrations according to different intentions and solving the IRL problem for each cluster. However, unlike HIQL, these frameworks do not accommodate the scenario where intentions might alternate within the same episode.
@@
In contrast, several approaches formulate MI-IRL problem as finding the maximum-likelihood partition of each trajectory, where each segment was generated from different locally consistent reward functions. To solve the problem, [DR12, MH12] extended the Bayesian IRL frameworks [RA07] to multi-intention scenario by introducing a Dirichlet process prior on different reward functions. The resulting Dirichlet process mixture model (DPM) can be solved via probabilistic inference, e.g., with Markov chain Monte Carlo [Nea00]. These algorithms do not require the number of intentions as a hyperparameter due to the nonparametric nature of DPMs. However, because of the nonparametric setting, DPMs inadequately model the temporal persistence of states and often create redundant states and rapidly switch among them, which eventually leads to poor interpretability of the results [Fox09]. This problem was addressed by Surana & Srivastava [SS14], where they extended the DPM to a sticky hierarchical Dirichlet process hidden Markov model. Nevertheless, solving the IRL problem as Bayesian inference is still computationally intensive, which could be intractable even for moderately sized finite-state IRL problems. Under the same problem formulation, Nguyen et al. [NLJ15] proposed a probabilistic graphical model, generalizing the algorithm from [BMSL11]. This approach avoids the computationally intensive Bayesian inference problem, but is restricted to the case of linearly-solvable environments. Our parametric HIQL algorithm also adapts the ‘trajectory segmentation + IRL’ problem formulation, but on the other hand, avoids the aforementioned limitations present in the other approaches. Particularly, we propose that in the parametric setup, manually specifying the number of intentions allows the researchers to decide the level of abstraction on intentions according to their expertise and scientific interest, i.e., whether to focus on the fine-grained difference between intentions (cf. §5.3).
@@
The state-of-the-art framework, DIRL [AJP22], can be considered as an extension of the maximum entropy IRL algorithm [ZMB+08, ZBD10] to non-stationary rewards, where the reward function is assumed to be parameterized as a time-varying linear combination of a small number of non-linear spatial reward maps with Gaussian random walk prior over weights. Under this assumption, the borderline between intentions in DIRL is blur, i.e., the intention transition is a continuous process, instead of a step-like procedure. However, such continuous time-varying reward assumed by DIRL is contradicted by the findings from [ARS+22], where it was proposed that humans and animals switch between multiple discrete strategies during decision-making. Although in theory, one can try to approximate such step-like intention transition dynamics with DIRL by assigning a larger variance $\sigma$ to the Gaussian to make the random walk less smooth, we show empirically (§5.1) that there are only minor influence from this hyperparameter $\sigma$ on the learnt reward, indicating that it would be challenging for DIRL to capture such step-like intention transition function during natural decision-making behavior. Instead, HIQL assumes that the intention transition dynamics follows a Markov process, which is intrinsically discrete and aligns better to the discrete switching intentions than DIRL (cf. §5.1 and §5.2).
@@ 3 Background
A Markov decision process (MDP) can be denoted by a tuple $\langle\mathcal{S},\mathcal{A},P,r,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ denotes the state- and action-space, respectively; The function $P\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition function with $P(s,a,s^{\prime})=\mathbf{P}(s^{\prime}\mid s,a)$ and $\mathbf{1}^{T}P(s,a,\cdot)=1$; The function $r\colon\mathcal{S}\times\mathcal{A}\to\mathbf{R}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor. Additionally, the function $\pi\colon\mathcal{S}\times\mathcal{A}\to[0,1]$ with $\pi(s,a)=\mathbf{P}(a\mid s)$ and $\mathbf{1}^{T}\pi(s,\cdot)=1$ is used to represent the policy according to which actions are selected in the MDP.
@@
Given the set of expert demonstrations $\mathcal{D}$ in an MDP, where each trajectory $\xi\in\mathcal{D}$ is denoted with a sequence of state-action pairs: $\xi=\left\{\left(s_{1},a_{1}\right),\ldots,\left(s_{n},a_{n}\right)\right\}$, the IRL problem consists in determining a reward function $r$ that best explains the observed expert behavior in the form of demonstrated trajectories. Unfortunately, IRL is generally ill-posed because infinitely many reward functions are consistent with the expert’s observed behavior. To resolve this issue, various approaches have been proposed (see [AD21] for a detailed review). Particularly, Kalweit et al. [KHWB20] formulated the IRL problem as a maximum likelihood estimation (MLE) problem:
@@
$\begin{array}[]{ll}\mbox{maximize}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\log\mathbf{P}\left(\xi\mid\pi_{r}\right)\right]\\ \mbox{subject to}&\mbox{$\pi_{r}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{$Q(s,a)=r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$},\end{array}$ (1)
@@
where $r$ is the optimization variable and $\mathcal{D}$ is the problem data. By assuming that the expert policy satisfies a Boltzmann distribution, if the environment model $P$ is known, problem (1) can be solved in closed-from via least squares [KHWB20], leading to the model-based inverse action-value iteration (IAVI) algorithm. On the other hand, when $P$ is unknown, the problem can be addressed via stochastic approximation, giving rise to the model-free inverse Q-learning (IQL) algorithm [KHWB20]. This approach needs to solve the MDP underlying the demonstrated behavior only once, leading to a speedup of up to several orders of magnitude compared to the popular maximum entropy IRL algorithm [ZMB+08] and some of its variants. In addition, it can accommodate arbitrary non-linear reward functions.
@@ 4 Hierarchical inverse Q-learning
Each expert demonstration is generated according to the Boltzmann optimal policy under one of the reward functions in a $K$-dimensional finite set =$\mathcal{R}=\left\{r_{1},\ldots,r_{K}\right\}$.
@@
The probability that one demonstration is generated under reward function $r\in\mathcal{R}$ is controlled by a Markov chain with initial state distribution $\Pi$ and transition matrix $\Lambda$, where the vector $\Pi$ and each row of the transition matrix $\Lambda_{i:}$, $i=1,\ldots,K$ represent some probability distribution over the set $\mathcal{R}$, i.e., $\Pi\in\Delta^{K-1}$ and $\Lambda_{i:}\in\Delta^{K-1}$, where $\Delta^{K-1}=\left\{x\in\mathbf{R}^{K}\mid x\succeq 0,\mathbf{1}^{T}x=1\right\}$ is a $K$-dimensional probability simplex.
@@
The resulting hierarchical decision process of an expert following these assumptions from the perspective of a researcher is depicted graphically in Figure 1, where the first two rows represent the MDP, and the last row represents the Markov chain for intention transition dynamics. Solving IRL problems on such decision network with parameter space $\Theta=\{\Pi,\Lambda,\mathcal{R}\}$ consists in determining 1) a set of reward functions, and 2) the reward function index for each demonstration that best jointly explain the observed expert behavior. An expectation-maximization (EM) algorithm can be devised to iteratively learn $\Theta$. For convenience, we introduce $\eta=\left\{z_{1},\ldots,z_{n}\right\}$ to be the predicted sequence of reward function indexes for trajectory $\xi\in\mathcal{D}$. Then each iteration of the EM process can be written as an MLE problem:
$\begin{array}[]{ll}\mbox{maximize}&J\left(\Theta^{+}\mid\Theta\right)=\mathbf{E}_{\xi\sim\mathcal{D},\eta}\left[\log\mathbf{P}\left(\xi,\eta\mid\Theta^{+}\right)\right],\end{array}$ (2)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Pi^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathbf{P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]\\ \mbox{\rm subject to}&\Pi^{+}\succeq 0,\ \mathbf{1}^{T}\Pi^{+}=1,\end{array}$ (3)
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Lambda^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]\\ \mbox{\rm subject to}&\Lambda^{+}_{i:}\succeq 0,\ \mathbf{1}^{T}\Lambda^{+}_{i:}=1,\quad i=1,\ldots,K,\end{array}$ (4)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $r_{i}^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{t=0}^{n}\mathbf{P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]\\ \mbox{\rm subject to}&\mbox{\rm$\pi_{r_{i}^{+}}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{\rm$Q(s,a)=r_{i}^{+}(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}.\end{array}$ (5)
@@
See Appendix A. ∎
In practice, to evaluate the objective functions of (3), (4) and (5), the Baum-Welch algorithm [BP66] can be applied to obtain the required posterior probabilities $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$ and $\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$ (cf. Appendix B). Then according to the Gibbs’ inequality, the optimum of (3) and (4) are achieved by
$\Pi^{+}_{i}=\mathbf{E}_{\xi\sim\mathcal{D}}[\mathbf{P}(z_{0}=i\mid\xi,\Theta)],\quad i=1,\ldots,K,$ (6)
@@
$\Lambda^{+}_{ij}=\frac{\mathbf{E}_{\xi\sim\mathcal{D},t}\left[\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\right]}{\mathbf{E}_{\xi\sim\mathcal{D},t}\left[\mathbf{P}(z_{t-1}=i\mid\xi,\Theta)\right]},\quad i=1,\ldots,K,\quad j=1,\ldots,K.$ (7)
To solve problem (5), note that it has the same structure as (1), thus it can be addressed by the class of IQL algorithms [KHWB20] with slight adaptations. Specifically, a weight of $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$ will be introduced to the $t$th demonstration during sampling.
Assembled, we call the above process of solving multi-intention IRL problems the hierarchical inverse Q-learning (HIQL) algorithm. The corresponding pseudo-code is listed in Algorithm 1.
Algorithm 1 Hierarchical inverse Q-learning.
given expert demonstrations $\mathcal{D}$ and reward set cardinality $K$.
1:initialize $\Pi$, $\Lambda$, $r_{1},\ldots,r_{K}$.
2:repeat
3:     Calculating the posterior probabilities $\mathbf{P}(z_{0}=i\mid\xi,\Theta)$, $\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$, and $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$ for each demonstration using the Baum-Welch algorithm.
4:     Update $\Pi$ and $\Lambda$ according to (6) and (7).
5:     for all $r\in\mathcal{R}$ do
6:         Update $r$ by solving problem (5) via IQL with each demonstration weighted by $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$.
7:     end for
8:until stopping criterion is satisfied.

### v4

@@ 1 Introduction
Characterizing decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019). Prior research has formulated a variety of mathematical behavioral models across diverse tasks (Ashwood et al., 2022b; Beron et al., 2022), including generalized linear models and models based on reinforcement learning. These forward models facilitate the understanding and comparison of decision-making strategies employed by both human and animal subjects. Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022). Forward models require an empirically defined reward function that guides subjects optimizing their behavior during decision-making. However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks (Alyahyay et al., 2023; Rosenberg et al., 2021). Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was maximizing its long-term return. Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as a valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al. (2018); Kwon et al. (2020); Alyahyay et al. (2023).
@@
Classic IRL methods seek to identify a single, fixed reward function for a specific scenario. In contrast, Ashwood et al. (2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity. Under this assumption, they proposed the dynamic inverse reinforcement learning (DIRL) framework, which parametrizes the animal’s reward function as a smoothly time-varying linear combination of a small number of spatial reward maps, which are referred to as ‘goal maps’. By assuming the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time. This innovative framework achieved state-of-the-art performance in animal behavior prediction. Nevertheless, demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al. (2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
@@
To address this requirement, we propose the novel class of hierarchical inverse Q-learning (HIQL) algorithms, which extend the fixed-reward inverse Q-learning (IQL) framework from Kalweit et al. (2020) to solve multi-intention IRL problems. Based on the assumption that the intention transition dynamics follows a Markov process, HIQL integrates an expectation-maximization (EM) approach to first divide expert trajectories into multiple intention segments in an unsupervised manner, and then solve the IRL problem independently for each segment. We then compare the performance of HIQL and the state-of-the-art model DIRL (Ashwood et al., 2022a) in a simulated gridworld environment and a dataset of trajectories from real mice navigating a $127$-node-labyrinth (Rosenberg et al., 2021). The results show that our HIQL algorithm outperforms DIRL in behavior prediction in both benchmarks, and produces interpretable reward functions. Finally, we applied HIQL in a real mice decision-making dataset from a dynamic two-armed bandit task (De La Crompe et al., 2023), and mathematically characterized exploitation and exploration behavior of animals during value-based decision-making.
@@ 2 Related work
Various approaches have been introduced to address multi-intention IRL problems. Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches allow for learning from multiple agents with distinct reward functions. Based on the assumption that the expert sticks to the same intention through one episode, these methods iterate between clustering expert demonstrations according to different intentions and solving the IRL problem for each cluster. However, unlike HIQL, these frameworks do not accommodate the scenario where intentions might alternate within the same episode.
@@
In contrast, several approaches formulate MI-IRL problem as finding the maximum-likelihood partition of each trajectory, where each segment was generated from different locally consistent reward functions. To solve the problem, Dimitrakakis & Rothkopf (2012); Michini & How (2012) extended the Bayesian IRL frameworks (Ramachandran & Amir, 2007) to multi-intention scenario by introducing a Dirichlet process prior on different reward functions. The resulting Dirichlet process mixture model (DPM) can be solved via probabilistic inference, e.g., with Markov chain Monte Carlo (Neal, 2000). These algorithms do not require the number of intentions as a hyperparameter due to the nonparametric nature of DPMs. However, because of the nonparametric setting, DPMs inadequately model the temporal persistence of states and often create redundant states and rapidly switch among them, which eventually leads to poor interpretability of the results (Fox, 2009). This problem was addressed by Surana & Srivastava (2014), where they extended the DPM to a sticky hierarchical Dirichlet process hidden Markov model. Nevertheless, solving the IRL problem as Bayesian inference is still computationally intensive, which could be intractable even for moderately sized finite-state IRL problems. Under the same problem formulation, Nguyen et al. (2015) proposed a probabilistic graphical model, generalizing the algorithm from Babes et al. (2011). This approach avoids the computationally intensive Bayesian inference problem, but is restricted to the case of linearly-solvable environments. Our parametric HIQL algorithm also adapts the ‘trajectory segmentation + IRL’ problem formulation, but on the other hand, avoids the aforementioned limitations present in the other approaches. Particularly, we propose that in the parametric setup, manually specifying the number of intentions allows the researchers to decide the level of abstraction on intentions according to their expertise and scientific interest, i.e., whether to focus on the fine-grained difference between intentions (cf. §5.3).
@@
The state-of-the-art framework, DIRL (Ashwood et al., 2022a), can be considered as an extension of the maximum entropy IRL algorithm (Ziebart et al., 2008; 2010) to non-stationary rewards, where the reward function is assumed to be parameterized as a time-varying linear combination of a small number of non-linear spatial reward maps with Gaussian random walk prior over weights. Under this assumption, the borderline between intentions in DIRL is blur, i.e., the intention transition is a continuous process, instead of a step-like procedure. However, such continuous time-varying reward assumed by DIRL is contradicted by the findings from Ashwood et al. (2022b), where it was proposed that humans and animals switch between multiple discrete strategies during decision-making. Although in theory, one can try to approximate such step-like intention transition dynamics with DIRL by assigning a larger variance $\sigma$ to the Gaussian to make the random walk less smooth, we show empirically (§5.1) that there are only minor influence from this hyperparameter $\sigma$ on the learnt reward, indicating that it would be challenging for DIRL to capture such step-like intention transition function during natural decision-making behavior. Instead, HIQL assumes that the intention transition dynamics follows a Markov process, which is intrinsically discrete and aligns better to the discrete switching intentions than DIRL (cf. §5.1 and §5.2).
@@ 3 Background
A Markov decision process (MDP) can be denoted by a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state- and action-space, respectively; The function $P\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition function with $P(s,a,s^{\prime})=\mathop{\bf P}(s^{\prime}\mid s,a)$ and $\mathbf{1}^{T}P(s,a,\cdot)=1$; The function $r\colon\mathcal{S}\times\mathcal{A}\to\mathbf{R}$ defines the reward function, and $\gamma\in[0,1]$ denotes the discount factor. Additionally, the function $\pi\colon\mathcal{S}\times\mathcal{A}\to[0,1]$ with $\pi(s,a)=\mathop{\bf P}(a\mid s)$ and $\mathbf{1}^{T}\pi(s,\cdot)=1$ is used to represent the policy according to which actions are selected in the MDP.
@@
Given the set of expert demonstrations $\mathcal{D}$ in an MDP, where each trajectory $\xi\in\mathcal{D}$ is denoted with a sequence of state-action pairs: $\xi=\left\{\left(s_{0},a_{0}\right),\ldots,\left(s_{n},a_{n}\right)\right\}$, the IRL problem consists in determining a reward function $r$ that best explains the observed expert behavior in the form of demonstrated trajectories. Note that both the word ‘demonstration’ and ‘trajectory’ will be used subsequently. The former refers to one action execution and state transition of the expert, while the latter corresponds to a sequence of actions and state transitions for an entire episode. Unfortunately, IRL is generally ill-posed because infinitely many reward functions are consistent with the expert’s observed behavior. To resolve this issue, various approaches have been proposed (see Arora & Doshi (2021) for a detailed review). Particularly, Kalweit et al. (2020) formulated the IRL problem as a maximum likelihood estimation (MLE) problem:
@@
$\begin{array}[]{ll}\mbox{maximize}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\log\mathop{\bf P}\left(\xi\mid\pi_{r}\right)\right]\\ \mbox{subject to}&\mbox{$\pi_{r}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{$Q(s,a)=r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$},\end{array}$ (1)
where $r$ is the optimization variable and $\mathcal{D}$ is the problem data. The objective of problem (1) is maximized when the expert policy $\pi^{\mathrm{E}}(s,a)$ is equivalent to $\pi_{r}(s,a)$. Hence,
$\pi^{\mathrm{E}}(s,a)=\frac{\exp(Q(s,a))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(s,a^{\prime}))}\implies Q(s,a)=Q(s,b)+\log(\pi^{\mathrm{E}}(s,a))-\log(\pi^{\mathrm{E}}(s,b)),$ (2)
for all actions $a\in\mathcal{A}$ and $b\in\mathcal{A}_{-a}$ where $\mathcal{A}_{-a}=\mathcal{A}\backslash\{a\}$. Using the Bellman optimality equation in (2), the immediate reward of action $a$ in state $s$ can be expressed by the immediate reward of some other action $b\in\mathcal{A}_{-a}$, the respective log-probabilities and future action-values (Kalweit et al., 2020):
$r(s,a)=\mu(s,a)+\frac{1}{\mathop{\bf card}(\mathcal{A}_{-a})}\sum_{b\in\mathcal{A}_{-a}}\left(r(s,b)-\mu(s,b)\right),$ (3)
where $\mathop{\bf card}(\mathcal{A}_{-a})$ denotes the cardinality of $\mathcal{A}_{-a}$, and
$\mu(s,a)=\log(\pi^{\mathrm{E}}(s,a))-\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime}).$ (4)
Let $\mathop{\bf card}(\mathcal{A})=m$, for each state $s\in\mathcal{S}$, formulating (3) for all actions $a\in\mathcal{A}$ results in a system of linear equations about $r(s,\cdot)$:
$\begin{bmatrix}1&-\frac{1}{m-1}&\cdots&-\frac{1}{m-1}\\ -\frac{1}{m-1}&1&\cdots&-\frac{1}{m-1}\\ \vdots&\vdots&\ddots&\vdots\\ -\frac{1}{m-1}&-\frac{1}{m-1}&\cdots&1\\ \end{bmatrix}\begin{bmatrix}r(s,a_{1})\\ r(s,a_{2})\\ \vdots\\ r(s,a_{m})\\ \end{bmatrix}=\begin{bmatrix}\mu(s,a_{1})-\frac{1}{m-1}\sum_{b\in\mathcal{A}_{-a_{1}}}\mu(s,b)\\ \mu(s,a_{2})-\frac{1}{m-1}\sum_{b\in\mathcal{A}_{-a_{2}}}\mu(s,b)\\ \vdots\\ \mu(s,a_{m})-\frac{1}{m-1}\sum_{b\in\mathcal{A}_{-a_{m}}}\mu(s,b)\\ \end{bmatrix}.$ (5)
@@
Thus the unknown reward function $r$ can be obtained by solving (5) for all $s\in\mathcal{S}$ via least squares (Kalweit et al., 2020). This approach is called the inverse action-value iteration (IAVI) (Algorithm 1), a model-based algorithm that solves the IRL problem analytically in closed-form. When the transition dynamics $P$ is unknown, stochastic approximation can be applied to (3) and (4) such that $r$ can be obtained in a sampling-based manner (Kalweit et al., 2020). The resulting inverse Q-learning (IQL) algorithm (Algorithm 2) is a model-free extension of IAVI. The class of IQL algorithms solve the MDP underlying the demonstrated behavior only once, leading to a speedup of up to several orders of magnitude compared to the popular maximum entropy IRL algorithm (Ziebart et al., 2008) and some of its variants. In addition, it can accommodate arbitrary non-linear reward functions.
@@ 4 Hierarchical inverse Q-learning
Each expert demonstration is generated according to the Boltzmann optimal policy under one of the reward functions in a $K$-dimensional finite set =$\mathcal{R}=\left\{r_{1},\ldots,r_{K}\right\}$, with each corresponding to one specific intention.
@@
The probability that one demonstration is generated under reward function $r\in\mathcal{R}$ is controlled by a Markov chain with initial state distribution $\Pi$ and transition matrix $\Lambda$, where the vector $\Pi$ and each row of the transition matrix $\Lambda_{i:}$, $i=1,\ldots,K$ represent some probability distribution over the set $\mathcal{R}$, i.e., $\Pi\in\Delta^{K-1}$ and $\Lambda_{i:}\in\Delta^{K-1}$, where $\Delta^{K-1}=\left\{x\in\mathbf{R}^{K}\mid x\succeq 0,\mathbf{1}^{T}x=1\right\}$ is a $(K-1)$-dimensional probability simplex.
@@
The resulting hierarchical decision process of an expert following these assumptions from the perspective of a researcher is depicted graphically in Figure 1, where the first two rows represent the MDP, and the last row represents the Markov chain for intention transition dynamics. Solving IRL problems on such decision network with parameter space $\Theta=\{\Pi,\Lambda,\mathcal{R}\}$ consists in determining 1) a set of reward functions, and 2) the reward function index for each demonstration that best jointly explain the observed expert behavior. An expectation-maximization (EM) algorithm can be devised to iteratively learn $\Theta$. For convenience, we introduce $\eta=\left\{z_{0},\ldots,z_{n}\right\}$ to be the predicted sequence of reward function indexes for trajectory $\xi\in\mathcal{D}$. Then each iteration of the EM process can be written as an MLE problem where the likelihood is obtained by marginalizing out the sequence of predicted reward function indexes $\eta$:
$\begin{array}[]{ll}\mbox{maximize}&J\left(\Theta^{+}\mid\Theta\right)=\mathop{\bf E}_{\xi\sim\mathcal{D},\eta}\left[\log\mathop{\bf P}\left(\xi,\eta\mid\Theta^{+}\right)\right],\end{array}$ (6)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Pi^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]\\ \mbox{\rm subject to}&\Pi^{+}\succeq 0,\ \mathbf{1}^{T}\Pi^{+}=1,\end{array}$ (7)
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Lambda^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]\\ \mbox{\rm subject to}&\Lambda^{+}_{i:}\succeq 0,\ \mathbf{1}^{T}\Lambda^{+}_{i:}=1,\quad i=1,\ldots,K,\end{array}$ (8)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $r_{i}^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{t=0}^{n}\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]\\ \mbox{\rm subject to}&\mbox{\rm$\pi_{r_{i}^{+}}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{\rm$Q(s,a)=r_{i}^{+}(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}.\end{array}$ (9)
@@
See Appendix A.1. ∎
In practice, to evaluate the objective functions of (7), (8) and (9), the Baum-Welch algorithm (Baum & Petrie, 1966) can be applied to obtain the required posterior probabilities $\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$ and $\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$ (cf. Appendix A.2). Then according to the Gibbs’ inequality, the optimum of (7) and (8) are achieved by
$\Pi^{+}_{i}=\mathop{\bf E}_{\xi\sim\mathcal{D}}[\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)],\quad i=1,\ldots,K,$ (10)
@@
$\Lambda^{+}_{ij}=\frac{\mathop{\bf E}_{\xi\sim\mathcal{D},t}\left[\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\right]}{\mathop{\bf E}_{\xi\sim\mathcal{D},t}\left[\mathop{\bf P}(z_{t-1}=i\mid\xi,\Theta)\right]},\quad i=1,\ldots,K,\quad j=1,\ldots,K.$ (11)
To solve problem (9), note that it has the same structure as (1), thus it can be addressed by the class of IQL algorithms (Algorithm 1 and 2) with slight adaptations. Specifically, for each $i=1,\ldots,K$, a subset of expert demonstrations from $\mathcal{D}$ will be sampled w.r.t. the posterior probability $\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$, then either Algorithm 1 or Algorithm 2 (depending on whether the transition dynamics $P$ is available) will be applied to this subset of demonstrations so that $r_{i}^{+}$ can be obtained. Assembled, we call the above process of solving multi-intention IRL problems the hierarchical inverse Q-learning (HIQL) algorithm. The corresponding pseudo-code is listed in Algorithm 3.